OpenStudy (anonymous):

Write the following real numbers in order form least to greatest: 1/624, 0.008, 5/972, 0.43%

OpenStudy (anonymous):

0.008 =8/1000 0.43% = 43/100 can youn finish it alone?

OpenStudy (anonymous):

yep thx!